OpenStudy (anonymous):

Write the equation in slope-intercept form: 2x + 3y = 18

OpenStudy (anonymous):

2x + 3y = 18 3y = 18 - 2x = - 2x + 18 y = (- 2x + 18) / 03 y = (- 2 / 3) x + 06

jimthompson5910 (jim_thompson5910):

\[\Large 2x + 3y = 18\] \[\Large 3y = 18-2x\] \[\Large 3y = -2x+18\] \[\Large y = \frac{-2x+18}{3}\] \[\Large y = -\frac{2}{3}x+\frac{18}{3}\] \[\Large y = -\frac{2}{3}x+6\]